### v1

# Learning Individual Reproductive Behavior from Aggregate Fertility Rates via Neural Posterior Estimation

Abstract
While age-specific fertility rates (ASFRs) provide the most extensive record of reproductive change, their aggregate nature masks the underlying behavioral mechanisms that ultimately drive fertility trends. To recover these mechanisms, we develop a likelihood-free Bayesian framework that couples an individual-level model of the reproductive process with Sequential Neural Posterior Estimation (SNPE). This allows us to infer eight behavioral and biological parameters from just two aggregate series: ASFRs and the age-profile of planned versus unplanned births. Applied to U.S. National Survey of Family Growth cohorts and to Demographic and Health Survey cohorts from Colombia, the Dominican Republic, and Peru, the method reproduces observed fertility schedules and, critically, predicts out-of-sample micro-level distributions of age at first sex, inter-birth intervals, and family-size ideals, none of which inform the estimation step. Because the fitted model yields complete synthetic life histories, it enables behaviorally explicit population forecasts and supports the construction of demographic digital twins.
Keywords— simulation-based inference, individual-level modelling, age-specific fertility rates, reproductive behavior, microsimulation
1 Introduction
Globally declining fertility rates are driving a profound demographic transformation, presenting first-order societal challenges worldwide and underscoring the critical need for robust models to understand and anticipate future fertility trends. Age-specific fertility rates (ASFRs) —births per woman at each age— remain the workhorse input for these tasks because the underlying birth counts are collected continuously in almost every country’s vital-statistics system. ASFRs series therefore provide unparalleled temporal and geographic coverage.
Yet ASFRs are only marginal aggregates. They reveal when births occur but not why. A long tradition of parametric models has been developed to describe ASFRs curves, smooth observed schedules, extrapolate fitted parameters, or convert projected total fertility rates into age patterns for cohort-component population forecasts (Coale and Trussell, 1974; Brass, 1974; Schmertmann, 2003; Chandola et al., 1999). While existing approaches can excel at capturing aggregate ASFRs shapes, their parameters are typically macro-level abstractions. As such, they offer limited or elusive insight into the individual behavioural mechanisms driving fertility decisions (Hoem and Hoem, 1989). Consequently, forecasts based on these models often cannot explicitly account for how evolving individual choices regarding the timing of parenthood, contraceptive practice, and birth-spacing and stopping decisions shape future fertility trends.
This disconnect is striking in light of the explosion of micro-level fertility research based on rich individual data sources. Studies using birth histories, time-use diaries, and linked administrative records have uncovered nuanced relationships between education, labor market trajectories, union stability, and childbearing. Yet virtually none of that behavioural insight is incorporated into fertility projection models. The challenge of understanding contemporary fertility lies in bridging this micro-macro divide.
This paper proposes a step toward this goal. We introduce a demographically interpretable individual-level stochastic model that captures key features of modern reproductive behavior, including stopping, spacing, and contraceptive failure. Estimating such a richly parameterised model from coarse data is a non-trivial inverse problem: many combinations of micro parameters can generate similar ASFRs curves. We show that adding the age profile of planned versus unplanned births to the set of ASFRs, provides enough information to identify the key parameters of this model. We fit the model to U.S. National Survey of Family Growth cohorts and to Demographic and Health Survey cohorts from Colombia, the Dominican Republic, and Peru, representing four settings that span markedly different fertility regimes.
Using Sequential Neural Posterior Estimation (SNPE), a likelihood-free method that learns to approximate posterior parameter distributions from simulator outputs, we recover full posterior distributions of behaviourally meaningful parameters. Posterior predictive checks further confirm that the inferred parameters enable the simulator to accurately reproduce the observed aggregate fertility and unplanned fertility rates. Most significantly, we conduct out-of-sample validation, demonstrating that the parameters inferred solely from these aggregate population rates can successfully predict distributions of individual-level characteristics, such as the empirical distributions of age at first sexual intercourse, desired family sizes reported in surveys, and realized birth intervals. This capacity to predict micro-level outcomes not directly used in the estimation process provides strong evidence for the model’s structural validity and the inference procedure’s power.
By establishing a demographically interpretable and statistically robust link between individual reproductive behaviors and population-level fertility outcomes, our framework offers a promising path towards more nuanced understanding and potentially more accurate and behaviorally grounded fertility forecasts.
1.1 Previous Work
Efforts to model the reproductive process mechanistically have a long and distinguished tradition in demography. Pioneered by early work on fecundability (Gini, 1924) and significantly advanced by work on ’natural fertility’ (Henry, 1953), this foundation led to a first wave of mathematical and early simulation models of the reproductive process (Brass, 1958; Singh, 1963; Ridley and Sheps, 1966; Potter, 1972; Bongaarts, 1977, among others). These models sought to capture the biological and behavioral drivers of family building with increasing realism.
However, this vibrant tradition of individual-level modeling did not continue to evolve within demography at the same pace or in the same manner as in other scientific fields. While empirical fertility research using micro-data flourished (Xie, 2000), the further development of these process-based simulators saw less sustained momentum. More critically, this stream of work remained largely disconnected from the concurrent revolution in statistical methodology for simulation-based inference. Techniques like Approximate Bayesian Computation (ABC) or other simulation-based inference (SBI) approaches that became influential in fields like ecology and genetics (Beaumont, 2010; Hartig et al., 2011), were not systematically integrated in demography. This methodological lag created a persistent gap in formally linking mechanistic understanding to empirical observations in fertility research.
Our own previous work showed that such a bridge is feasible in natural-fertility contexts —historical or religious communities without deliberate birth control— where a handful of biological parameters govern the reproductive process and can be inferred from ASFRs alone (Ciganda and Todd, 2024). The present paper tackles the far more ambitious, and policy-relevant task, of modelling contemporary fertility, where stopping, spacing, and imperfect contraception introduce complex decision making and heterogeneity in reproductive dynamics.
1.2 Organization of the Article
The remainder of this paper is structured as follows: Section 2 describes our individual-level computational model of the reproductive process. Section 3 details the data sources and the construction of ASFRs and Age-specific unplanned fertility rates (ASUFRs). Section 4 outlines the SNPE methodology. The Results section presents our findings on parameter identifiability via cross-validation, model fit through posterior predictive checks, and the out-of-sample validation against micro-level data. Finally, we discuss the implications of our findings, limitations, and avenues for future research in the Discussion.
2 Model
The transition from a natural to a modern fertility regime in human populations is marked by the emergence of stopping and spacing behaviors (deciding whether and when to have children), representing a shift from a reproductive process largely determined by physiological factors to one shaped by individual preferences. As a result, fertility patterns become more diverse, with family formation increasingly tailored to individual or couple-level choices rather than biological constraints alone.
Building on these insights, our model represents the dynamics of reproductive behavior in contemporary populations by placing fertility intentions and desires at its core. Specifically, it conceptualizes childbearing decisions as the interplay between physiological capacity, access to contraception, and individuals’ subjective preferences, such as desired family size and timing.
We employ an individual-level, discrete-time simulation model. A cohort of $N$ women is simulated from the age of potential sexual activity until the end of their reproductive capacity, with their state updated in monthly time steps.
The model begins by assigning each woman key characteristics that will shape her reproductive life. These include two critical ages: the age at sexual initiation, when she first becomes exposed to the risk of childbearing, and the age at intentional reproduction, representing the point at which she actively starts trying to conceive. Because these two ages often differ, there is typically a period during which a woman is sexually active but not yet intending to have children, an interval associated with the risk of unplanned pregnancies.
Additionally, each woman is assigned a desired family size, reflecting the number of children she ultimately hopes to have, and a desired number of months between births, which influences her birth spacing behavior.
In each monthly step, a woman’s probability of conception is evaluated. This probability depends on her current age, her current parity relative to her desired family size, whether she has reached her age at intentional reproduction, and, if she has had prior births, whether her desired birth spacing interval has elapsed.
Specifically, when a woman is past her age at intentional reproduction, has not yet reached her desired family size, and is past her desired birth spacing interval (if applicable), we assume she is actively trying to conceive and uses no contraception. Conversely, if she has not yet reached her age at intentional reproduction, or has already achieved her desired family size, or is within her desired spacing interval, she is assumed to be using contraception. This results in a reduced, but non-zero, probability of conception, accounting for imperfect contraceptive effectiveness and the possibility of unplanned pregnancies throughout the reproductive lifespan.
Whenever a conception occurs, the woman is considered non-susceptible for nine months of pregnancy, followed by an additional three-month period of postpartum amenorrhea, during which her probability of another conception is zero. After giving birth, a further delay is introduced to represent the woman’s desired number of months between births, mirroring real-life behaviors around birth spacing.
2.1 Implementation
The dynamics described above are implemented by assigning the following characteristics to each woman $i$ and defining the processes that govern her reproductive journey:
  1. 1.
    Age at Sexual Initiation ($X_{i}$): $X_{i}$ (in months) is drawn from a lognormal distribution. This distribution is parameterized by an estimated mean age at sexual initiation $\mu_{s}$ (in years) and an estimated standard deviation $\sigma_{s}$ (in years). These are converted to the lognormal distribution’s underlying parameters (mean of logarithms $\mu_{\ln,s}$ and standard deviation of logarithms $\sigma_{\ln,s}$) using the standard formulas: $\mu_{\ln,s}=\ln(\mu_{s}^{2}/\sqrt{\mu_{s}^{2}+\sigma_{s}^{2}})$ and $\sigma_{\ln,s}=\sqrt{\ln(1+\sigma_{s}^{2}/\mu_{s}^{2})}$. Ages are then drawn and converted to months.
  2. 2.
    Age at Intentional Reproduction ($R_{i}$): $R_{i}$ (in months) is also drawn from a lognormal distribution. The mean of this distribution is set to $\mu_{r}=\mu_{s}+\delta_{r}$, where $\delta_{r}$ (in years) is an estimated model parameter. The standard deviation for the age at intentional reproduction, $\sigma_{r}$, is fixed at 2.6 years. The parameters $\mu_{r}$ and $\sigma_{r}$ (in years) are used to derive the underlying lognormal parameters $\mu_{\ln,r}$ and $\sigma_{\ln,r}$ using the same conversion formulas.
  3. 3.
    Desired Family Size ($D_{i}$): $D_{i}$ is drawn from a Weibull distribution, parameterized by an estimated mean desired family size $\mu_{d}$ and a fixed standard deviation $\sigma_{d}=1.25$ children. These are used to approximate the Weibull shape parameter $\alpha_{d}=(\sigma_{d}/\mu_{d})^{-1.086}$ and scale parameter $\lambda_{d}=\mu_{d}/\Gamma(1+1/\alpha_{d})$. Samples are rounded to the nearest integer.
  4. 4.
    Desired Birth Spacing ($B_{i}$): $B_{i}$ (in months) is drawn from a lognormal distribution, parameterized by an estimated mean desired spacing $\mu_{b}$ (in months) and standard deviation $\sigma_{b}=27$ months. The underlying lognormal parameters $\mu_{\ln,b}$ and $\sigma_{\ln,b}$ are derived similarly.
We fix the dispersion parameters ($\sigma_{r},\sigma_{d},\sigma_{b}$) at plausible, empirically-informed values because, while contributing to overall individual heterogeneity, these second-order moments were found to exert a less direct or distinguishable influence on the aggregate ASFRs and ASUFR patterns compared to parameters governing central tendencies or the core dynamics of conception and contraception. Furthermore, robustly identifying these specific dispersion parameters solely from the available aggregate fertility schedules proved challenging in preliminary analyses. Fixing them allows the inferential procedure to concentrate more effectively on the primary behavioral and physiological parameters that are both more influential at the aggregate level and better identified by the data, without excessively constraining the model’s ability to capture essential individual-level variability.
The age-specific monthly probability of conception (fecundability), $\phi(x)$ at age $x$ (in years), is a crucial component of the model. It is well established that fecundability typically increases from puberty, reaches a peak during the early adult years, and progressively declines thereafter towards menopause (Bendel and Hua, 1978; Larsen and Yan, 2000; Weinstein et al., 1990). To capture this characteristic age pattern, we model $\phi(x)$ over a reproductive window from $x_{min}=10$ to $x_{max}=50$ years. Age $x$ is first rescaled to $x_{s}=(x-x_{min})/(x_{max}-x_{min})$, so $x_{s}\in[0,1]$. Instead of using a standard polynomial basis, we employ Bernstein basis polynomials, which provide flexibility and ensure the curve is well-behaved at the boundaries of the reproductive window. Specifically, $\phi(x)$ is modeled as a linear combination of two Bernstein basis polynomials of degree 3:
$\phi(x)=\beta_{1}[3x_{s}(1-x_{s})^{2}]+\beta_{2}[3x_{s}^{2}(1-x_{s})]$
The terms $3x_{s}(1-x_{s})^{2}$ and $3x_{s}^{2}(1-x_{s})$ correspond to the two intermediate basis functions ($B_{1,3}(x_{s})$ and $B_{2,3}(x_{s})$) of a Bernstein basis of degree 3. The exclusion of the first and last basis functions from the model inherently forces the curve $\phi(x)$ to be zero at the start ($x_{s}=0$) and end ($x_{s}=1$) of the reproductive window. The parameters $\beta_{1}$ and $\beta_{2}$ are estimated and control the level and shape of the fecundability curve, allowing for a variety of empirically plausible age-fecundability profiles.
The actual probability of conception in a given month depends not only on this baseline age-specific fecundability but also on a woman’s reproductive intentions and contraceptive use. As outlined earlier, if a woman $i$ with current parity $k_{i}$ is actively trying to conceive (i.e., her current age is $\geq R_{i}$, her parity $k_{i}<D_{i}$, and sufficient time $B_{i}$ has elapsed since her last birth), her probability of conception is $\phi(x)$.
Conversely, if a woman $i$ with current parity $k_{i}$ is sexually active but not actively trying to conceive, she is assumed to use contraception. Her monthly probability of conception is then $\kappa_{eff}\phi(x)$. The parameter $\kappa\in(0,1)$ is an estimated baseline probability of contraceptive failure. To model an increased contraceptive vigilance once desired parity ($D_{i}$) is met, the effective failure rate $\kappa_{eff}$ is defined as $\kappa$ if her current parity $k_{i}<D_{i}$, and is reduced to $\kappa^{2}$ if $k_{i}\geq D_{i}$. This formulation reflects intensified efforts to avoid further births, substantially reducing the risk of an unplanned pregnancy (assuming $\kappa<1$) after a woman reaches her desired family size.
Upon simulating complete reproductive trajectories, we derive aggregate age-specific fertility rates (ASFRs). However, relying solely on ASFRs for parameter inference may be insufficient given the depth of behavioral detail incorporated in our model. Without distinguishing births that align with stated intentions from those that are mistimed or unwanted, different parameter configurations (e.g., high desired family size with effective contraception versus lower desired family size with poor contraception) could yield deceptively similar ASFRs, creating critical identifiability challenges. Therefore, alongside ASFRs, a key output of our model is a vector of age-specific unplanned fertility rates (ASUFRs). These ASUFRs, detailed further in Section 3, exclusively take into account births classified as mistimed or unwanted, providing essential information to better constrain the model’s parameters and narrow the range of configurations consistent with the observed data.
Table 1 summarizes the key estimated model parameters.
Parameter Description
$\mu_{s}$ Mean age at sexual initiation (in years), modeled with a lognormal distribution.
$\sigma_{s}$ Standard deviation for the distribution of ages at sexual initiation.
$\delta_{r}$ Mean gap (in years) between sexual initiation and intentional reproduction.
$\mu_{d}$ Mean desired family size, modeled using a Weibull distribution.
$\kappa$ Probability of contraceptive failure.
$\mu_{b}$ Mean desired birth spacing (in months), modeled with a lognormal distribution.
$\beta_{1}$ Defines the peak of the age-specific fecundability curve.
$\beta_{2}$ Defines the decline of the age-specific fecundability curve.
Table 1: Key model parameters and their descriptions.
3 Data
Our analysis draws on two large, nationally representative sources. For countries outside the United States, we use the Demographic and Health Surveys (DHS), a leading initiative of the United States Agency for International Development (USAID) that has conducted over 300 surveys in more than 90 low-and middle-income countries. Each DHS round collects complete retrospective birth histories, allowing the reconstruction of every woman’s fertility trajectory. Data for the United States are drawn from several early cycles of the National Survey of Family Growth (NSFG), specifically the 1972, 1976, 1982, and 1988 rounds. These surveys, conducted by the National Center for Health Statistics (NCHS), are designed to provide detailed national data on family life, marriage, divorce, contraception, and importantly for this study, complete pregnancy and birth histories. We utilize these data as compiled and harmonized by the Integrated Fertility Survey Series (IFSS), applying IFSS sampling weights to generate population-level estimates. For both DHS and NSFG data, to achieve adequate sample sizes while ensuring cohorts of women were exposed to broadly similar social conditions, we pool available survey rounds and organize respondents into 10-year birth cohorts.
A key consideration in selecting empirical settings was to align them with the behavioral logic embedded in our model. Thus, we narrowed the DHS universe to countries where adolescent childbearing is primarily unintended. This was operationalized by calculating, for each potential cohort, the proportion of births to women under 18 that were declared as unplanned, retaining only those populations where this share exceeded one-half. This selection criterion identifies populations where, for a significant fraction of teenagers, sexual activity and intentional reproduction are decoupled—a precondition for our model’s assumption that early fertility is largely viewed as undesirable. It also serves to exclude contexts where prevailing norms, such as marriage-centered childbearing from a young age, would imply a data-generating process markedly different from the one our model specifies.
Because DHS sample sizes and the number of survey rounds differ widely, we implemented a further filter based on cohort size. We selected DHS countries where the chosen cohort included a minimum of 4,000 women, a threshold established to guarantee reasonably precise estimates of both age-specific total and unplanned fertility rates across the entire reproductive span. This resulted in the inclusion of three DHS countries: Colombia, Peru, and the Dominican Republic. Together with the United States, these four nations constitute our study sample, offering a valuable cross-section of different fertility patterns and socio-economic contexts that allows for a robust examination of our model’s performance across diverse demographic and social settings.
We compute the conventional age-specific fertility rate (ASFRs) by dividing the number of births at each age by the number of women in the cohort at that age. To construct the age-specific unplanned fertility rate (ASUFR) we combine each birth’s intention status with the respondent’s ideal family size. In the DHS, questions about birth intention (i.e., whether a birth was “wanted then,” “wanted later,” or “not wanted”) are asked only for events in the 3–5 years preceding the interview. We therefore estimate, for every maternal age, the share of below-ideal births classed as unplanned in that recent window and extrapolate it to earlier events, while always counting births that push parity above the ideal as unwanted. In the NSFG, by contrast, intention status is recorded for all births, so we classify each event directly from the survey response and apply the same parity rule for above-ideal births. By coupling direct reports where available with an indirect parity check where necessary, our strategy mitigates recall bias and ex-post rationalization, which are known to severely affect unplanned fertility estimates obtained from the intention questions alone (Casterline and El-Zeini, 2007).
4 Simulation-Based Inference
Inferring parameters for complex process-based models, such as our individual-level simulation of reproductive trajectories, is often challenging due to intractable likelihood functions. Neural Posterior Estimation (NPE) offers a powerful simulation-based inference (SBI) approach that circumvents these difficulties by directly learning an approximation to the posterior distribution $p(\theta\mid x)$ from simulated data pairs $(\theta,x)$. In practice, NPE proceeds by drawing parameters $\theta$ from a prior distribution ${p}(\theta)$, using the simulator to generate the corresponding data $x\sim p(x\mid\theta)$, and then training a flexible neural density estimator, $q_{\phi}(\theta\mid x)$ (with network parameters $\phi$), on these simulated pairs to approximate the true posterior.
While basic NPE is effective, our study employs a powerful sequential variant known as Automatic Posterior Transformation (APT), sometimes referred to as SNPE-C (Greenberg et al., 2019). Like other sequential neural posterior estimation (SNPE) methods, APT enhances inferential efficiency by drawing parameter samples from an iteratively updated proposal distribution, $\tilde{p}(\theta)$, rather than solely from the fixed prior. The key challenge in this approach is to correct for the use of a proposal distribution, as naively training on the resulting samples would lead to an incorrect posterior estimate. APT solves this by analytically transforming the posterior estimate itself. For each simulated data point drawn from a proposal $\tilde{p}_{r}(\theta)$, it re-weights the target density using the known ratio $\tilde{p}_{r}(\theta)/p(\theta)$. The network is then trained by minimizing the loss with respect to this transformed "proposal posterior" (see line 10 of Algorithm 1). A key advantage of this formulation is that because the loss is valid for any proposal, the network can be trained on the cumulative set of all simulations generated across all previous rounds. This procedure allows the network to learn the true posterior without relying on the potentially unstable importance weights used by other methods.
A key strength of the APT framework is its flexibility, as it supports a wide range of density estimators, including mixture-density networks and various types of normalizing flows. In this study, we employ a Neural Spline Flow (NSF), an expressive and powerful density estimator based on monotonic rational-quadratic splines (Durkan et al., 2019).
This adaptive refinement progressively concentrates simulation effort on regions of high posterior probability, leading to a more accurate posterior approximation with fewer overall simulations. SNPE and related neural posterior estimation techniques have found successful application in diverse scientific domains, including neuroscience (Gonçalves et al., 2020; Groschner et al., 2022; Deistler et al., 2022), astrophysics (Dax et al., 2025), cognitive science (von Krause et al., 2022), and exoplanet searches (Vasist et al., 2023), among others.
Algorithm 1 summarizes the complete APT procedure employed in this study, which was implemented using the sbi Python package (Boelts et al., 2024).
Input: Simulator with implicit likelihood $p(x|\theta)$, observed data $x_{o}$, prior $p(\theta)$, conditional density family $q_{F(x,\phi)}(\theta)$, number of rounds $R$, simulations per round $N$.
Output: The final posterior approximation $q_{F(x_{o},\phi)}(\theta)$.
Set initial proposal distribution $\tilde{p}_{1}(\theta)\leftarrow p(\theta)$;
for $r=1$ to $R$ do
       for $j=1$ to $N$ do
               Sample $\theta_{r,j}\sim\tilde{p}_{r}(\theta)$;
               Simulate $x_{r,j}\sim p(x|\theta_{r,j})$;
        end for
       Update network weights $\phi$ by minimizing the loss over all collected data $\{(\theta_{i,j},x_{i,j})\}$ from rounds $i=1...r$:
       $\phi\leftarrow\operatorname*{arg\,min}_{\phi}\sum_{i=1}^{r}\sum_{j=1}^{N}-\log\tilde{q}_{x_{i,j},\phi}(\theta_{i,j})$;
[…]
       Set the next proposal to the current posterior estimate for the observed data $x_{o}$:
       $\tilde{p}_{r+1}(\theta)\leftarrow q_{F(x_{o},\phi)}(\theta)$;
end for
Algorithm 1 Automatic Posterior Transformation (APT)
The code and data used in this study, along with detailed instructions for replication, can be found on GitHub at github.com/dciganda/comfert_regulated. The repository is currently private but access can be granted upon request to the corresponding author.
5 Results
5.1 Cross-Validation for Parameter Identifiability
Simulation-Based Inference (SBI) typically relies on summarizing potentially high-dimensional data into lower-dimensional, sufficiently informative statistics. This can be done either through explicit, pre-defined summaries or via learned embeddings. The ability to work with a low-dimensional yet descriptive representation of the data is often essential for making parameter inference tractable, especially for complex simulators.
When working with aggregate data, as in our study, the dimensionality reduction is an inherent feature of the observation process. For instance, vital statistics or administrative registers implicitly aggregate thousands of individual fertility trajectories into relatively low-dimensional summaries like age-specific fertility rates (ASFRs). We thus inherit a pre-summarized dataset, limiting our ability to engineer or learn optimal summary statistics directly from the underlying micro-level process.
A critical question then arises: do these aggregate data carry sufficient information to uniquely identify the parameters of our detailed individual-level model? Different combinations of micro-parameters (e.g., those governing desired family size, timing of intentional reproduction, or contraceptive failure) could potentially yield very similar aggregate fertility schedules, posing identifiability challenges.
To rigorously assess the degree to which our model parameters can be recovered from aggregate data, we conduct extensive validation experiments using simulated data. These experiments evaluate parameter recovery under three increasingly informative inference setups, varying the summary statistics provided to the SNPE algorithm and the strength of prior information for key parameters:
  1. 1.
    ASFRs-only: Parameters are drawn from weakly informative priors. The SNPE algorithm is conditioned using only simulated age-specific fertility rates (ASFRs) as summary statistics.
  2. 2.
    ASFRs + ASUFR: Priors remain weakly informative. The summary statistics are augmented to include both simulated ASFRs and age-specific unplanned fertility rates (ASUFRs), providing direct information on the planned versus unplanned nature of births.
  3. 3.
    ASFRs + ASUFR + informative priors on $\mu_{d}$ and $\kappa$: Summary statistics are ASFRs and ASUFRs, as in Scenario 2. However, we now incorporate plausible external information by using narrower, more informative priors for the mean desired family size ($\mu_{d}$) and the baseline contraceptive failure rate ($\kappa$). All other parameters retain their weakly informative priors.
For each scenario, our validation procedure involves repeatedly generating datasets with known parameters and then attempting to recover these parameters using our full inference pipeline. Specifically, we repeat the following steps for $I$ iterations:
  1. 1.
    Generate Ground-Truth Data: In each iteration $i$ (for $i=1,\dots,I$), we draw a distinct "true" parameter vector, denoted $\theta^{*(i)}$, from the prior distribution $p(\theta)$ defined for the given scenario. Using our individual-level simulation model, we then generate a corresponding vector of aggregate summary statistics (e.g., ASFRs and ASUFRs), denoted $x_{0}^{(i)}$, which serves as the pseudo-observed data for this iteration. The pair $(\theta^{*(i)},x_{0}^{(i)})$ constitutes a ground-truth dataset where the data-generating parameters are known.
  2. 2.
    Perform Full Posterior Inference: For each pseudo-observed dataset $x_{0}^{(i)}$, we independently run our complete Sequential Neural Posterior Estimation (SNPE) procedure. This involves training a neural density estimator over $R$ rounds (in our application, $R=8$), with $N_{sim}$ new simulations (in our application, $N_{sim}=1000$) drawn from the evolving proposal distribution in each round, to obtain an approximation of the posterior distribution $q_{\phi}(\theta\mid x_{0}^{(i)})$.
  3. 3.
    Evaluate Parameter Recovery: From each resulting posterior approximation $q_{\phi}(\theta\mid x_{0}^{(i)})$, we compute a point estimate for the parameters, typically the posterior mean, denoted $\hat{\theta}^{(i)}$. We then assess the accuracy of recovery by comparing this estimate $\hat{\theta}^{(i)}$ with the known true generating parameters $\theta^{*(i)}$.
By performing these $I$ independent inference runs, we obtain a distribution of estimation errors (e.g., $\hat{\theta}^{(i)}-\theta^{*(i)}$) for each model parameter under each of the three informational scenarios. We then summarize these error distributions using the root-mean-squared error (RMSE). Lower RMSE values indicate better parameter recovery and allow for a direct comparison of how effectively each informational setup constrains the model parameters.
Figure 1 reports the normalized RMSE for all parameters under our three scenarios. The ASFRs-only scenario shows higher RMSE across parameters, indicating that ASFRs alone provide weaker identifiability for several model components. The incorporation of information on the planned versus unplanned nature of births via ASUFRs (Scenario 2) contributes to significantly reducing these errors, particularly for parameters governing the timing of reproductive intentions ($\delta_{r}$) and the rate of contraceptive failure ($\kappa$). Finally, the use of auxiliary information to inform the priors for $\mu_{d}$ and $\kappa$ (Scenario 3) yields further improvements, highlighting how complementary data can refine parameter identifiability.
Scenario 1 Scenario 2 Scenario 3 Normalized RMSE
$\mu_{s}$
$\sigma_{s}$
$\delta_{r}$
$\mu_{d}$
$\kappa$
$\mu_{b}$
$\beta_{1}$
$\beta_{2}$
Figure 1: Cross-validation of eight model parameters over 50 folds. Columns 1–3 display scatterplots of true parameter values (horizontal axis) versus posterior median estimates (vertical axis) for Scenarios 1, 2, and 3, with the dashed $45^{\circ}$ reference line indicating perfect recovery. Column 4 reports the scenario-specific root-mean-squared error.
5.2 Model Validation
The cross-validation experiments (Section 5.1) confirmed the capacity of our SBI framework to recover key micro-level parameters from aggregate ASFRs and ASUFR data. To further scrutinize the model’s performance, we conduct two additional forms of validation. The first, posterior predictive checks, assesses the model’s ability to replicate the aggregate data patterns used for inference, thus evaluating its internal consistency and goodness-of-fit. The second, and more demanding, validation stage examines the model’s capacity to predict out-of-sample individual-level characteristics of the observed cohorts, such as distributions of desired family size, age at first sexual intercourse, and birth interval durations, which were not used during the parameter estimation process.
For both validation exercises, we focus on parameter posteriors obtained under Scenario 2 (ASFRs + ASUFR with weakly informative priors). This scenario is chosen for its parsimonious data requirements (relying only on the age-specific share of unplanned births alongside total ASFRs) while having demonstrated strong parameter recovery in the cross-validation exercises.
5.2.1 Posterior Predictive Checks for Aggregate Rates
To perform posterior predictive checks, we drew 5,000 samples from the joint posterior distribution of the model parameters for each of the four countries (United States, Colombia, Dominican Republic, and Peru). For each parameter sample, we simulated a full cohort of individual reproductive histories using our model and then recomputed the aggregate ASFRs and ASUFRs series.
Figure 2 presents the empirically observed ASFRs and ASUFRs curves against their corresponding posterior predictive distributions. In these plots, observed total ASFRs are marked by purple open circles and observed ASUFRs by purple open triangles. The red open circles and triangles indicate the mean of the 5,000 simulated ASFRs and ASUFRs schedules, respectively, while the shaded red areas depict the 97.5% posterior predictive intervals.
Across all four diverse demographic contexts, the model demonstrates a strong capacity to replicate the observed data. The posterior predictive means for both ASFRs and ASUFR closely track the empirical curves, capturing key features such as the location and height of peak fertility, the age pattern of decline, and the distinct levels and shapes of unplanned fertility. For instance, the distinct, lower peak of the ASUFRs in the US is well reproduced. For Colombia (Figure 2b), the Dominican Republic (Figure 2c), and Peru (Figure 2d), characterized by higher overall fertility, the model successfully replicates not only the higher ASFRs levels but also the specific age gradients of both total and unplanned births; in Colombia and Peru, however, there is a discernible tendency for the predictive mean to underestimate the peak of the ASFRs, though the observed data remain well within the prediction intervals. Notably, across all countries, almost all empirically observed points for both ASFRs and ASUFRs fall well within their respective 97.5% posterior predictive intervals, indicating that the model, conditioned on the inferred parameters, generates aggregate outcomes highly consistent with the observed data.
These results confirm that our framework, even with the relatively minimal data inputs of Scenario 2, can consistently and accurately reproduce the population-level fertility and unplanned fertility schedules it was intended to explain. This provides strong confidence in the internal validity of the estimated parameters before assessing their ability to predict out-of-sample, individual-level outcomes.
(a) United States
(b) Colombia
(c) Dominican Republic
(d) Peru
Figure 2: Posterior Predictive Checks for Age-Specific Fertility (ASFR) and Unplanned Fertility (ASUFR) Rates. Comparison of observed rates (purple open circles for ASFR, purple open triangles for ASUFR) with model-based posterior predictions for four countries: (a) United States, (b) Colombia, (c) Dominican Republic, and (d) Peru. Red open circles and triangles indicate the mean of 5,000 posterior predictive simulations for ASFR and ASUFR, respectively. Red shaded areas represent the 95% posterior predictive intervals. Parameters were estimated under Scenario 2 (ASFR + ASUFR with weakly informative priors).
5.2.2 Out-of-Sample Validation
To assess the model’s ability to generate realistic individual-level reproductive behaviors not directly used in parameter estimation, we use parameter values representative of the posterior distribution (specifically, the posterior mean obtained under Scenario 2, as described in Section 5.2) to drive new simulations. These simulations yield full synthetic reproductive histories. We then compare the distributions of key micro-level outcomes from these synthetic life courses, namely, age at first sexual activity, desired family size, and inter-birth intervals, with their empirical counterparts derived from the survey data (Figure 3). This side-by-side comparison provides a direct and demanding appraisal of the model’s realism in capturing individual-level reproductive patterns.
Overall fit.
Across the four countries and three micro-level indicators, the distributions generated from the model’s synthetic life courses closely approximate the empirical distributions observed in the survey data. This indicates that the parameter vectors inferred solely from aggregate ASFRs and ASUFRs schedules enable the reconstruction of distinct micro-behavioural dimensions within the observed cohorts, without any specific tuning towards these individual-level targets. Quantitatively, for instance, the Jensen–Shannon divergence (base 2) between simulated and observed distributions never exceeds 0.056 bits for any of the three distributions evaluated.
Age at first sex.
The simulated distributions reproduce the overall shape and cross-country ordering of sexual debut: the mean age is highest in the United States and Peru, slightly lower in Colombia, and lowest in the Dominican Republic, with model-based means only a few months off the survey values. Two departures stand out: the sharp heap of reports at exactly age 18 (likely a mix of social norms and recall heaping) is smoothed out by the model, and in the United States the simulated curve is displaced very slightly to the left of the observed one, suggesting debut a few months earlier than recorded. Even with these minor mismatches, the alignment is notable, considering that the entire distribution is inferred solely from aggregate fertility rates.
Desired family size.
Overall, the simulated preferences closely match the shape of the observed distributions, reproducing the main peak around two to three children and the gradual taper into larger family ideals. A clear deviation appears at exactly two children, where the empirical data show a sharp bump (a clear outcome of strong social norms) that the model’s simple two-parameter form inevitably smooths over. Beyond this focal point the fit remains close, with only a slightly thinner extreme right tail in the simulated curves.
Birth intervals.
The synthetic data recovers the principal spacing pattern: a broad peak around two to three years followed by a gradual decline towards longer gaps. The model undershoots the cluster of very short gaps and shows a slightly heavier tail beyond five years. Apart from these edges of the distribution, the match is close across the main mass of birth intervals.
Together, these out-of-sample checks demonstrate that, even when estimated solely from ASFRs and ASUFRs, our mechanistic model successfully reconstructs a rich array of individual reproductive behaviors.
Age at First Sex Desired Family Size Birth Interval
United States
Peru
Colombia
Dominican Rep.
Figure 3: Validation on micro-level outcomes. Purple bars denote the observed survey distributions, while red lines show the model-implied counterparts generated from posterior draws. Columns display key behavioral dimensions—age at first sex, desired family size, and birth interval—whereas rows correspond to each study cohort (United States, Peru, Colombia, Dominican Republic).
6 Discussion
This study has introduced a powerful likelihood-free Bayesian framework capable of inferring complex individual-level reproductive behaviors solely from aggregate population data. By integrating a demographically interpretable individual-level simulation model with Sequential Neural Posterior Estimation (SNPE), we demonstrated that eight core parameters (governing reproductive timing, quantum, and contraceptive efficacy) can be reliably estimated using only age-specific fertility rates and information on unplanned fertility. The most compelling validation of our approach lies in the model’s ability, using these inferred parameters, to accurately predict out-of-sample micro-level distributions for behaviors such as age at first sexual intercourse, desired family size, and birth intervals, none of which informed the original estimation. This establishes a new, statistically rigorous bridge between macro-level fertility observations and their underlying micro-behavioral drivers.
Methodologically, this work serves as a strong empirical demonstration of SNPE’s utility for parameter inference in interpretable, yet complex, social science simulation models where direct likelihood calculation is prohibitive. The successful estimation from just two aggregate series, and the critical role of ASUFRs in disambiguating behavioral parameters and enhancing identifiability, offer transferable insights for simulation-based inference across disciplines facing similar inverse problems. Substantively, our framework provides a novel toolkit for demographers, enabling the estimation of behaviorally meaningful parameters (e.g., mean desired family size, age at intentional reproduction, contraceptive failure rates) even in data-scarce contexts, thereby offering a mechanistic complement to traditional micro-data analyses.
This capability to infer interpretable behavioral parameters significantly enhances population forecasting, moving beyond traditional methods that often rely on extrapolating aggregate trends or employ macro-level parameters with limited behavioral grounding. Our approach enables the generation of ’behaviorally explicit population forecasts’ where future scenarios are constructed from hypothesized changes in these underlying, micro-level behavioral parameters. Crucially, by first distilling complex fertility dynamics into these more causally proximate behavioral components, our framework also facilitates a clearer understanding of how external covariates, such as education or economic conditions, shape reproductive outcomes. Machine learning models, for instance, may more effectively identify systematic relationships when predicting these interpretable behavioral parameters rather than attempting to directly predict highly aggregated and confounded ASFRs. This improved ability to model the determinants of specific behaviors not only leads to more robust and theoretically sound inputs for scenario-based population projections but also enriches the potential for constructing responsive ’demographic digital twins’ for policy exploration.
Despite these advances, the present framework has limitations that define important avenues for future development. The individual-level model, while behaviorally detailed, necessarily simplifies complex human realities; notably, select dispersion parameters for reproductive preferences were fixed to ensure precise inference of core behavioral drivers from aggregate data. The crucial ASUFRs are derived from survey data on birth intentions, which carry inherent measurement considerations (such as recall bias or post-hoc rationalization), even with our methodological safeguards. Furthermore, the model’s current conceptualization—predicated on individualized fertility intentions and the distinction between planned, mistimed, and unwanted births—is best suited to contemporary populations where reproduction is substantially decoupled from sexuality and where such articulable birth control decision-making is prevalent. This may limit its direct applicability to contexts with more natural fertility regimes or where reproductive agency and intentions are structured differently. A further practical constraint is that while ASFRs are nearly universally available, the ASUFR data necessary for the full inferential power demonstrated here are far less common, currently restricting the range of settings where the complete framework can be readily applied.
These limitations notwithstanding, this research opens up several exciting directions for future work, including strategies to broaden the framework’s applicability. Addressing the contextual specificity of the current model, future efforts could adapt its structure for diverse fertility regimes. For populations closer to natural fertility, the model could be simplified, focusing on a smaller set of parameters. In contexts where Western notions of "unplanned" births are less salient, research could explore identifying and incorporating alternative summary statistics that reflect local modes of fertility regulation.
Regarding the scarcity of ASUFR data, while this study underscores the urgent need for more systematic collection of information on birth intentions, several interim strategies could enhance the model’s utility. One avenue is the development of methods to generate plausible ASUFR estimates for data-sparse regions, potentially by leveraging covariate information (e.g., education, contraceptive prevalence) from data-rich settings or by "borrowing strength" from demographically similar populations. Alternatively, the framework could be employed using ASFRs alone but with significantly stronger, empirically-grounded priors on the behavioral parameters that ASUFRs primarily help identify; this would yield posterior distributions with greater uncertainty but could still offer valuable insights, particularly for sensitivity analyses under different assumptions about underlying birth intentions.
Appendix
The visualizations presented in Figures 4 to 7 present the outcomes of the Sequential Neural Posterior Estimation (SNPE) for the United States, Colombia, the Dominican Republic, and Peru cohorts respectively. These plots were generated under Scenario 2, where model parameters were inferred using both age-specific fertility rates (ASFRs) and age-specific unplanned fertility rates (ASUFRs) with weakly informative priors. The diagonal elements of each plot display the marginal posterior distributions for the eight estimated model parameters, while the off-diagonal elements show the bivariate joint posterior distributions, revealing correlations and interdependencies between parameter pairs.
Across all four cohorts, most marginal posterior distributions are unimodal and have been substantially updated from their initial weakly informative priors (see Figure 8 for a visual comparison of the posterior distributions against their respective priors). Parameters such as $\mu_{s}$ (mean age at sexual initiation), $\sigma_{s}$ (standard deviation of sexual initiation age), $\mu_{d}$ (mean desired family size), $\kappa$ (contraceptive failure probability), and the fecundability curve parameters ($\beta_{1}$ and $\beta_{2}$) generally exhibit well-defined and relatively concentrated posterior distributions. This implies that the ASFR and ASUFR data contain sufficient information to substantially constrain these posteriors, concentrating their probability mass within a relatively confined region of the parameter space.
While most posterior distributions are well defined, $\mu_{b}$ (mean desired birth spacing) and $\delta_{r}$ (mean gap between sexual initiation and intentional reproduction) exhibit more diffuse posterior distributions in certain national contexts, indicating varying degrees of constraint by the aggregate data. Specifically, $\delta_{r}$ presents a notably wide and less defined posterior for the U.S. cohort (see Figure 4). In contrast, $\mu_{b}$ tends to be the parameter showing greater uncertainty in Colombia (see Figure 5) and Peru (see Figure 7). The Dominican Republic stands out as a setting where both $\mu_{b}$ and $\delta_{r}$ appear relatively better constrained compared to the other cohorts (see Figure 6). Despite this variability, the model’s overall predictive performance, as assessed by the posterior predictive checks (Figure 2) and the out-of-sample validation of micro-level outcomes (Figure 3), remains strong.
The off-diagonal plots in Figures 47 are essential for understanding the interdependencies between parameters. Several notable patterns emerge from these bivariate distributions.
Among these, a strong and consistent negative correlation is observed between the fecundability parameters $\beta_{1}$ and $\beta_{2}$ across national contexts. This relationship is structurally expected, as these two parameters jointly define the age-specific fecundability curve. This interdependence implies that the overall shape of the fecundability curve is a well-identified feature, even if the exact individual values of $\beta_{1}$ and $\beta_{2}$ have some trade-off. Similarly, a general negative correlation is typically observed between $\mu_{s}$ (age at sexual initiation) and $\delta_{r}$ (gap to intentional reproduction) across cohorts. This inverse relationship is also expected, as these parameters jointly determine the overall timing of entry into intentional reproduction.
A discernible negative correlation is also generally present between the fecundability peak parameter $\beta_{1}$ and the contraceptive failure rate $\kappa$ in the cohorts examined. This particular interdependency, while not as directly apparent from the model’s structure as the others, likely reflects a trade-off in fitting our observed fertility aggregates. For instance, a scenario with higher intrinsic peak fecundability (larger $\beta_{1}$) could produce similar aggregate fertility outcomes to one with lower peak fecundability if compensated by a correspondingly lower contraceptive failure rate (smaller $\kappa$), particularly when trying to match the levels of unplanned births, which are influenced by contraceptive effectiveness. This suggests that the available aggregate data may not always fully disentangle the distinct contributions of these two parameters. Nonetheless, these trade-offs are relatively contained and do not significantly reduce the model’s overall predictive utility.
Figure 4: Posterior parameter distributions for the United States cohort (Scenario 2). The diagonal displays marginal posteriors for each of the eight model parameters, while off-diagonal plots show their bivariate relationships.
Figure 5: Posterior parameter distributions for the Colombian cohort (Scenario 2). The diagonal displays marginal posteriors for each of the eight model parameters, while off-diagonal plots show their bivariate relationships.
Figure 6: Posterior parameter distributions for the Dominican Republic cohort (Scenario 2). The diagonal displays marginal posteriors for each of the eight model parameters, while off-diagonal plots show their bivariate relationships.
Figure 7: Posterior parameter distributions for the Peruvian cohort (Scenario 2). The diagonal displays marginal posteriors for each of the eight model parameters, while off-diagonal plots show their bivariate relationships.
[…]
[figure]capposition=bottom
US CO DR PE
$\mu_{s}$
$\sigma_{s}$
$\delta_{r}$
$\mu_{d}$
Figure 8: Marginal posterior distributions of the eight key behavioral and biological model parameters for the United States, Colombia, Dominican Republic, and Peru, inferred from empirical data under Scenario 2 (Part 1 of 2).
US CO DR PE
$\kappa$
$\mu_{b}$
$\beta_{1}$
$\beta_{2}$
Figure 9: Marginal posterior distributions of the eight key behavioral and biological model parameters for the United States, Colombia, Dominican Republic, and Peru, inferred from empirical data under Scenario 2 (Part 2 of 2).